OpenStudy (anonymous):

a sound wave is modeled with the equation y = 1/4 cos 2(pi)/3 (then some weird zero with a line through it)... find the period and amplitude ...

OpenStudy (loser66):

I think you miss x after 2pi/3, right?

OpenStudy (anonymous):

no there is a weird circle after it

OpenStudy (anonymous):

u missed x after 2pi/3 right??

OpenStudy (anonymous):

yea u did

OpenStudy (anonymous):

there isn't an x on it, but I guess include the x if needed to solve it :/ i just want to know how to answer it

OpenStudy (anonymous):

oh im sorry im only 12

OpenStudy (loser66):

ok, if you don't have x, then it is a point on a graph, no amplitude, no period.

OpenStudy (loser66):

Suppose there is x there. I meant the function is \(y =(1/4) cos (2\pi/3) x\)

OpenStudy (loser66):

amplitude is 1/4

OpenStudy (anonymous):

yes what he said im onlt 12

OpenStudy (kropot72):

Is this the equation? \[\large \frac{1}{4}\cos \frac{2 \pi}{3} \theta\]

OpenStudy (anonymous):

yeah

OpenStudy (kropot72):

The period is given by: \[\large T=\frac{2 \pi}{\frac{2 \pi}{3}}=3\]

OpenStudy (anonymous):

thanks!

OpenStudy (kropot72):

You're welcome :)